# Using Low-Discrepancy Points for Data Compression in Machine Learning: An Experimental Comparison

### v1

@@ 1 Introduction
In many applications, the increase in the volume of data has led to several computational challenges, in particular in machine learning and the accompanying optimization problems.
Data reduction is a classical technique that reduces the size of a dataset while still preserving the most important information. Concepts and methods in this field include Core-Sets [6], support points [18] and random subsampling [17], to mention a few. However, low-discrepancy points and quasi Monte Carlo-techniques have received little attention in this context so far. We are only aware of the work [4], which was the starting point for our study.
@@ 2 Low-discrepancy point sets and regression
Let $\mathcal{P}=\left\{\boldsymbol{x}_{1},\ldots,\boldsymbol{x}_{N}\right\}\subset[0,1)^{s}$ and
@@
$D_{N}(\mathcal{P})=\sup_{\begin{subarray}{c}\boldsymbol{a},\boldsymbol{b}\in[0,1]^{s}\\ \boldsymbol{a}\leq\boldsymbol{b}\end{subarray}}\left|\frac{A([\boldsymbol{a},\boldsymbol{b}),N)}{N}-\lambda_{s}([\boldsymbol{a},\boldsymbol{b}))\right|$
@@
$D_{N}^{*}(\mathcal{P})=\sup_{\boldsymbol{a}\in[0,1]^{s}}\left|\frac{A([{0},\boldsymbol{a}),N)}{N}-\lambda_{s}([{0},\boldsymbol{a}))\right|.$
@@
Here $V(g)$ is the Hardy-Krause variation of $g$. See, e.g. Chapter 2 in [22] for these results and further information.
@@
Now using low-discrepancy points $P=\left\{\boldsymbol{z}_{1},\ldots,\boldsymbol{z}_{L}\right\}\subset[0,1)^{s}$ for the $dx$-integration and the data from $\mathcal{Y}$ for the $dy$-integration, one obtains an approximation of the form (1.3), i.e.
@@ 3.1 Quasi-Monte Carlo compression
For each $(t_{\alpha},m,s)$-net $P$ in base $b$ of order $\alpha$ and for each $\nu\leq m-t_{\alpha}$, $\boldsymbol{d}\in\mathbb{N}_{0}^{s}$ and $\boldsymbol{a}\in K_{\boldsymbol{d}}$, it can be shown that $\left|P_{\boldsymbol{a},\boldsymbol{d}}\right|=b^{m-\nu+q}$, where $P_{\boldsymbol{a},\boldsymbol{d}}$ denotes the set $P$ intersected with $I_{\boldsymbol{a},\boldsymbol{d}}$ and $\left|\boldsymbol{d}\right|=\nu-q$ for $q\in\{0,\ldots,\min(s-1,\nu)\}$. This is a consequence of Definition (3.1) and will be important for deriving the weights.
@@
Lemma (3.2) can be exploited to find an expression for the weights $\left\{W_{\mathcal{X},P,\nu,l}\right\}_{l=1}^{L}$. In the following, we assume that our data set is given by $\mathcal{X}=\left\{\boldsymbol{x}_{1},\ldots,\boldsymbol{x}_{N}\right\}$, our answers by $\mathcal{Y}=\left\{y_{1},\ldots,y_{N}\right\}$, our compressed data set by $P=\left\{\boldsymbol{z}_{1},\ldots,\boldsymbol{z}_{L}\right\}$, and that $\nu$ is fixed.
@@ 3.2 Supercompress method
In [14] the authors proposes another, more robust algorithm by taking the convex combination
@@ 3.3 QMC-Voronoi method
To construct the compressed data set for the QMC-Voronoi method, we start with a digital net. Instead of calculating complex weights, we form clusters by using the Voronoi diagram, see e.g. Section I.1 in [8], on the digital net. This implies that for each data point $\boldsymbol{x}$, the closest point $\boldsymbol{z}$ of the digital net $P$ is identified based on the Euclidean distance, and $\boldsymbol{x}$ is assigned to the Voronoi region of $\boldsymbol{z}$. Hence, $P$ represents the compressed set. For each compressed point $\boldsymbol{z}$, the corresponding answer $w$ is found by taking the average of the corresponding answers of the data points $\boldsymbol{x}$, which are in the Voronoi region of $\boldsymbol{z}$. Consequently, as for the supercompress method the error $\operatorname{err}(f_{\theta})$ can therefore be approximated by $\operatorname{app}_{L}^{clst}(f_{\theta})$. The above procedure is summarized in Algorithm (V), where we take $K=b^{m}$ points for the compressed data set.
@@ 4 Numerical results
Finally, we undertake a numerical comparison of the methods. We start by considering three fixed functions $f$ and compare $\operatorname{err}(f)$ with $\operatorname{app}_{L}^{avg}(f_{\theta})$ and $\operatorname{app}_{L}^{clst}(f_{\theta})$, respectively. Subsequently, we train a neural network using the MNIST dataset [16] and compare the accuracy achieved with the different compression methods. All implementations can be found in our Git [12].
@@ 4.1 Test functions
This function satisfies $\|f_{1}\|<\infty$ as well as $V(f_{1})<\infty$, but has no higher regularity. The second test function is given by
@@
This function is discontinuous, but still has finite Hardy-Krause variation, i.e. $V(f_{2})<\infty$. Finally, the so-called Zhou function is given by
@@
$\phi\left(\boldsymbol{x}\right)=\frac{1}{\left(2\pi\right)^{\frac{s}{2}}}\exp\left(-\frac{\left\|\boldsymbol{x}\right\|^{2}}{2}\right)$
@@
Note that these functions have different scales. For example, for $s=2$ the second function attains values between zero and $\max\{1,\exp(a_{1}u_{1})\}\max\{1,\exp(a_{2}u_{2})\}$, while the first one takes values between zero and one only. Finally, Zhou’s function contains values between zero and at most $7.9579$. In the following, we will use $u_{1}=\ldots=u_{s}=0.5$ and $a_{1}=\ldots=a_{s}=5$. For more information and illustrations of these functions, see [23].
@@
Table (1) illustrates the results for the dimension $s=2$. The top row expresses the average error of the supercompress algorithm, while the bottom row refers to that of the QMC-averaging method. It can be observed that the supercompress algorithm performs better for each function. Only for the discontinuous function $f_{2}$ and high compression rates, i.e. small $K$, both methods perform equally bad. Since $f_{2}([0,1]^{2})=[0,\exp(5)]$, and $\exp(5)\approx 148,4132$, the observed behavior might be a consequence of the lack of scaling. To analyze this, we repeat the procedure for $\gamma f_{2}/\|f_{2}\|_{\infty}^{2}$ with different scales $\gamma$. In Table (2) we see, that both methods yield a smaller error, if $\gamma$ is low. Thus, both methods behave similar for all test functions when the latter are properly scaled; the supercompress method seems to be always the superior choice. Interestingly, the regularity of the test functions does not seem to have a big influence on the performance of the QMC-averaging method. You can find a scaled version with $\gamma=1$ of all tables regarding the error analysis in the Appendix.
compression rate $N$ $s$ $K$ no scale scale 100 scale 10 scale 1
$9\%$ $3000$ $2$ $2^{8}$ $2,9351\cdot 10^{2}$ $1,3326\cdot 10^{0}$ $1,3326\cdot 10^{-1}$ $1,3326\cdot 10^{-2}$
$6,9552\cdot 10^{3}$ $3,1577\cdot 10^{1}$ $3,1577\cdot 10^{0}$ $3,1577\cdot 10^{-1}$
$17\%$ $3000$ $2$ $2^{9}$ $8,7618\cdot 10^{1}$ $3,9779\cdot 10^{-1}$ $3,9779\cdot 10^{-2}$ $3,9779\cdot 10^{-3}$
$6,4545\cdot 10^{3}$ $2,9304\cdot 10^{1}$ $2,9304\cdot 10^{0}$ $2,9304\cdot 10^{-1}$
@@
We also performed the same numerical experiment without adding noisy perturbations. The results are found to be consistent with those above, indicating that adding noise does not significantly influence the comparison.
The sensitivity of both methods in terms of the dimension is presented in Table (3). For $f_{1}$, the size of the problem does not affect the accuracy of our methods. However, it can be observed that there is a deterioration in accuracy for the other functions. In particular, the error increases significantly in higher dimensions when considering $f_{2}$.
@@ 4.2 Neural networks
Thus far, we have considered only uniformly random data and a rather simple connection between data points and responses. However, in general this is not the case. To illustrate this, we will now examine the MNIST data set, which can be found in [16]. This data set contains $6\cdot 10^{4}$ grey-scale images of handwritten numbers between 00 and $9$. Each image is represented by a matrix of dimensions $(28\times 28)$, with entries from 00 to $1$. The value of each entry indicates the darkness of the corresponding pixel. By concatenating all the pixels in a given image, we obtain the image of the handwritten number. Consequently, the set $\mathcal{X}$ contains the pixel sets***Here we identify $1$ as $1-{\tt eps}$, where ${\tt eps}$ is the machine accuracy. Thus, we can still work with the assumption $\mathcal{X}\subset[0,1)^{s}$., while the set $\mathcal{Y}$ contains the corresponding handwritten numbers. Figure 1 illustrates 16 examples of such pixel sets. In the case of an entry of 00, the pixel is depicted as black.
@@
A comparison of the performance of the compression algorithms with that of the uncompressed training is presented in Table (6). The table shows the accuracy, the required time (i.e. $12:34$ corresponds to 12 minutes and 34 seconds), and the final loss, with the exception of the QMC-averaging method, for which the complex input structure (weights instead of corresponding answers) does not allow a comparable loss value. The number of compressed points $L=K$ is varied for values in the set $\{2^{7},2^{9},2^{10},2^{11},2^{12}\}$, with the most interesting results selected.
@@
We observed a striking phenomenon about the weights. As previously stated in Section (3.1), the weights serve to indicate the relative importance of each point within the digital net. It is notable that for all compression rates, the weights associated with the first two points (of which one is always the zero vector) are considerably higher than those of the remaining points. Indeed, the latter are largely approximately zero. This implies that the majority of the points are located within the same elementary interval as the first two points for a significant number of partitions. Consequently, the QMC-averaging point set is not an accurate representation of the data set $\mathcal{X}$. This is primarily due to particular structure of $\mathcal{X}$. Since a substantial number of pixels are completely black (have a 0-entry), it is understandable that a large proportion of points is close to the 00-vector. Therefore, a representative set that is well distributed on the unit cube, is not a good choice for our data. In fact, this behavior can be even more drastically observed when attempting to use the QMC-Voronoi algorithm to train the network. It is noteworthy that all points from the set $\mathcal{X}$ are assigned to only two points (in fact, the first two points) in the set $P$. Consequently, these points are the only points for which a corresponding answer exists and, therefore, the only points to train the neural network with. Given that two points are insufficient for training a neural network, this method could not be included in Table (6).
The reason why $K$-means clustering is not performing well may be attributed to the clustering based on the $\boldsymbol{x}$-space. This causes many clusters to contain points with different corresponding answers. This is not a problem for a cluster like $\{2,2,2,2,2,2,2,4,4\}$, since the rounded cluster mean response is still $2$. Though, clusters such as $\{2,2,2,2,4,4,4,4\}$ correspond to an answer of $3$, which is not even an answer within the cluster. Therefore incorporating the $y$-space into the clustering method appears to be a logical approach. This could also explain why the robust version performs less effectively in this instance. It incorporates the $\boldsymbol{x}$-space into the loss function instead of just the $y$-space. As a result, for this discrete $y$-space problem, a clustering approach based on the output space, performs the best. One potential solution to this issue is to employ a different rounding method. Instead of taking the average of the answers, we could select the answer that appears the most frequently. However, this does not address the fundamental challenge that methods primarily using the $\boldsymbol{x}$-space to cluster can produce clusters with a multitude of corresponding answers.
@@
The supercompress is the only method that predicts every number (with a certain degree of accuracy). The robust version fails to predict zeros. Consequently, it overpredicts other numbers, particularly high numbers such as $7$ and $8$. This is the primary reason for the lower accuracy compared to the normal supercompress. The $K$-means clustering method drastically overestimates $2$ and $3$ to be the correct numbers. This systematic error highlights the ineffectiveness of this method in this context. The most unsatisfactory result is produced by QMC-averaging, which predicts only one number. This is not surprising, given that only two points have notable weights and therefore dominate the loss function. The reason for the absence of certain classes in the prediction can be explained by Table (7), which shows the distribution of the corresponding answers among the original and compressed data sets. It should be noted that due to rounding, the percentages may not sum to $100$. Additionally, there is a greater concentration of mass around high numbers. This is a consequence of the fact that many clusters contain two points. If the average value has a $0.5$ part, the answer is rounded up, for example, the answer corresponding to the set $\{8,9\}$ is 9.

### v2

@@ 1 Introduction
Data reduction is a classical technique that reduces the size of a dataset while still preserving the most important information. Concepts and methods in this field include Core-Sets [6], support points [19] and random subsampling [18], to mention a few. However, low-discrepancy points and quasi Monte Carlo-techniques seem to have received little attention in this context so far. They were used to create training data points for learning surrogate models in urban traffic by [26] and also similarly in [17, 24] for a variety of applications, including several types of partial differential equations. Dick and Feischl [4] proceeded differently, namely by using low-discrepancy points to compress known data for the training of neural networks. This work was the starting point of our study.
@@ 2 Low-discrepancy point sets and regression
We start by introducing the concept of low-discrepancy point sets and some corresponding results. See, e.g. Chapter 2 in [23] for further information. Let $\mathcal{P}=\left\{\boldsymbol{x}_{1},\ldots,\boldsymbol{x}_{N}\right\}\subset[0,1)^{s}$ and
@@
$D_{N}(\mathcal{P})=\sup_{\begin{subarray}{c}\boldsymbol{a},\boldsymbol{b}\in[0,1]^{s}\\ \boldsymbol{a}\leq\boldsymbol{b}\end{subarray}}\left|\frac{A([\boldsymbol{a},\boldsymbol{b}),\mathcal{P})}{N}-\lambda_{s}([\boldsymbol{a},\boldsymbol{b}))\right|$
@@
$D_{N}^{*}(\mathcal{P})=\sup_{\boldsymbol{a}\in[0,1]^{s}}\left|\frac{A([{0},\boldsymbol{a}),\mathcal{P})}{N}-\lambda_{s}([{0},\boldsymbol{a}))\right|.$
@@
Here and in the following $V(g)$ is the Hardy-Krause variation of $g$.
@@
Now using low-discrepancy points as the compressed data points $P=\left\{\boldsymbol{z}_{1},\ldots,\boldsymbol{z}_{L}\right\}\subset[0,1)^{s}$ for the $dx$-integration and the data from $\mathcal{Y}$ for the $dy$-integration, one obtains an approximation of the form (1.3), i.e.
@@ 3.1 Quasi-Monte Carlo compression
For each $(t_{\alpha},m,s)$-net $P$ in base $b$ of order $\alpha$ and for each $\nu\leq m-t_{\alpha}$, $\boldsymbol{d}\in\mathbb{N}_{0}^{s}$ and $\boldsymbol{a}\in K_{\boldsymbol{d}}$, it can be shown that $\left|P_{\boldsymbol{a},\boldsymbol{d}}\right|=b^{m-\nu+q}$, where $P_{\boldsymbol{a},\boldsymbol{d}}$ denotes the set $P$ intersected with $I_{\boldsymbol{a},\boldsymbol{d}}$ and $\left|\boldsymbol{d}\right|=\nu-q$ for $q\in\{0,\ldots,\min(s-1,\nu)\}$. As mentioned in Chapter 2 of [4], this is a consequence of Definition (3.1) and will be important for deriving the weights.
@@
Lemma (3.2) can be exploited to find an expression for the weights $\left\{W_{\mathcal{X},P,\nu,l}\right\}_{l=1}^{L}$. In the following, we assume that our data set is given by $\mathcal{X}=\left\{\boldsymbol{x}_{1},\ldots,\boldsymbol{x}_{N}\right\}$, our responses by $\mathcal{Y}=\left\{y_{1},\ldots,y_{N}\right\}$, our compressed data set by $P=\left\{\boldsymbol{z}_{1},\ldots,\boldsymbol{z}_{L}\right\}$, and that $\nu$ is fixed.
@@ 3.2 Supercompress method
In [14] the authors propose another, more robust algorithm by taking the convex combination
@@
We will refer to this more robust algorithm, which optimizes $\widetilde{L}$, as robust supercompress.
@@ 3.3 QMC-Voronoi method
To construct the compressed data set for the QMC-Voronoi method, we start with a digital net. Instead of calculating complex weights, we form clusters by using the Voronoi diagram, see e.g. Section I.1 in [8], on the digital net. This implies that for each data point $\boldsymbol{x}$, the closest point $\boldsymbol{z}$ of the digital net $P$ is identified based on the Euclidean distance, and $\boldsymbol{x}$ is assigned to the Voronoi region of $\boldsymbol{z}$. Hence, $P$ represents the compressed set. For each compressed point $\boldsymbol{z}$, the corresponding response $w$ is found by taking the average of the corresponding responses of the data points $\boldsymbol{x}$, which are in the Voronoi region of $\boldsymbol{z}$. Consequently, as for the supercompress method the error $\operatorname{err}(f_{\theta})$ can therefore be approximated by $\operatorname{app}_{L}^{clst}(f_{\theta})$. The above procedure is summarized in Algorithm (V), where we take $K=b^{m}$ points for the compressed data set.
@@ 4 Numerical results
Finally, we undertake a numerical comparison of the methods. We start by considering three fixed functions $f$ and compare $\operatorname{err}(f)$ with $\operatorname{app}_{L}^{avg}(f_{\theta})$ and $\operatorname{app}_{L}^{clst}(f_{\theta})$, respectively. Subsequently, we train a neural network using the MNIST dataset [16] and compare the accuracy achieved with the different compression methods. All implementations can be found in our Git repository [12].
@@ 4.1 Test functions
This function satisfies $\|f_{1}\|<\infty$ as well as $V(f_{1})<\infty$, but has no higher regularity w.r.t. $\|\cdot\|_{p.\alpha}$ for $\alpha\geq 2$, due to the involved absolute values. The second test function is given by
@@
[…]
@@
$\phi\left(\boldsymbol{x}\right)=\frac{1}{\left(2\pi\right)^{\frac{s}{2}}}\exp\left(-\frac{\left\|\boldsymbol{x}\right\|^{2}}{2}\right).$
@@
Note that these functions have different scales. For example, for $s=2$ the second function attains values between zero and $\max\{1,\exp(a_{1}u_{1})\}\max\{1,\exp(a_{2}u_{2})\}$, while the first one takes values between zero and one only. Finally, Zhou’s function contains values between zero and at most $7.9579$. In the following we will choose $u_{1}=\ldots=u_{s}=0.5$ so that $f_{1}$ is symmetric around the center of the unit cube and the discontinuity of $f_{2}$ appears at the same position. Furthermore we set $a_{1}=\ldots=a_{s}=5$. Thus the maximum value of $f_{2}$ for $s=2$ is $(\exp(2.5))^{2}\approx 148.41$, which is much larger than that of $f_{1}$ and $f_{3}$. For more information and illustrations of these functions, see [25].
@@
Table (1) illustrates the results for the dimension $s=2$. The top row expresses the average error of the supercompress algorithm, while the bottom row refers to that of the QMC-averaging method. It can be observed that the supercompress algorithm performs better for each function. Only for the discontinuous function $f_{2}$ and high compression rates, i.e. small $K$, both methods perform equally bad. Since $f_{2}([0,1]^{2})=[0,\exp(5)]$, and $\exp(5)\approx 148,4132$, the observed behavior might be a consequence of the lack of scaling. To analyze this, we repeated the procedure for $\gamma f_{2}/\|f_{2}\|_{\infty}$ with different scales $\gamma$. In Table (2) we see, that both methods yield a smaller error for $\gamma=1$, which corresponds to standardized test functions. Thus, both methods behave similar for all test functions when the latter are properly scaled; the supercompress method seems to be always the superior choice. Interestingly, the regularity of the test functions does not seem to have a big influence on the performance of the QMC-averaging method. You can find a scaled version with $\gamma=1$ of all tables regarding the error analysis in the Appendix (Table (8), Table (9), Table (10)).
compression rate $N$ $s$ $K$ no scale scale 1
$9\%$ $3000$ $2$ $2^{8}$ $2,9351\cdot 10^{2}$ $1,3326\cdot 10^{-2}$
$6,9552\cdot 10^{3}$ $3,1577\cdot 10^{-1}$
$17\%$ $3000$ $2$ $2^{9}$ $8,7618\cdot 10^{1}$ $3,9779\cdot 10^{-3}$
$6,4545\cdot 10^{3}$ $2,9304\cdot 10^{-1}$
@@
We also performed the same numerical experiment without adding noisy perturbations. The results are found to be consistent with those above, indicating that adding noise does not significantly influence the comparison. In fact, as long as the noise level of the data is reasonably low, the effects of the above analysis remain the same.
The sensitivity of both methods in terms of the dimension is presented in Table (3). For $f_{1}$, the size of the problem does not affect the accuracy of our methods. However, it can be observed that there is a deterioration in accuracy for the other functions. In particular, the error increases significantly in higher dimensions when considering $f_{2}$. Note that for the largest dimension $s=10$, QMC-averaging leads to smaller errors for $f_{1}$ and $f_{3}$ than supercompress. We do not have a mathematical explanation at hand and did not explore this phenomenon further due to the very high running time of QMC-averaging for larger $s$; see also Table (5).
@@ 4.2 Neural networks
Thus far, we have considered only uniformly random data and a rather simple connection between data points and responses. However, in general this is not the case. To illustrate this, we will now examine the MNIST data set, which can be found in [16]. This data set contains $6\cdot 10^{4}$ grey-scale images of handwritten numbers between 00 and $9$. Each image is represented by a matrix of dimensions $(28\times 28)$, with entries from 00 to $1$. The value of each entry indicates the darkness of the corresponding pixel. By concatenating all the pixels in a given image, we obtain the image of the handwritten number. Consequently, the set $\mathcal{X}$ contains the pixel sets***Here we identify $1$ as $1-{\tt eps}$, where ${\tt eps}$ is the machine accuracy. Thus, we can still work with the assumption $\mathcal{X}\subset[0,1)^{s}$., while the set $\mathcal{Y}$ contains the corresponding handwritten numbers. Figure (3) illustrates 16 examples of such pixel sets. In the case of an entry of 00, the pixel is depicted as black.
@@
A comparison of the performance of the compression algorithms with that of the uncompressed training is presented in Table (6). The table shows the accuracy, the required time (i.e. $12:34$ corresponds to 12 minutes and 34 seconds), and the final loss, with the exception of the QMC-averaging method, for which the complex input structure (weights instead of corresponding responses) does not allow a comparable loss value. The number of compressed points $L=K$ is varied for values in the set $\{2^{9},2^{10},2^{11},2^{12}\}$. We have dropped the results of QMC-averaging and $K$-means for $L=K=2^{12}$, due to no significant improvement in accuracy.
@@
We observed a striking phenomenon about the weights. As previously stated in Section (3.1), the weights serve to indicate the relative importance of each point within the digital net. It is notable that for all compression rates, the weights associated with the first two points (of which one is always the zero vector) are considerably higher than those of the remaining points. Indeed, the latter are largely approximately zero. This implies that the majority of the points are located within the same elementary interval as the first two points for a significant number of partitions. Consequently, the QMC-averaging point set is not an accurate representation of the data set $\mathcal{X}$. This is primarily due to particular structure of $\mathcal{X}$. Since a substantial number of pixels are completely black (have a 0-entry), it is understandable that a large proportion of points is close to the 00-vector. Therefore, a representative set that is well distributed on the unit cube, is not a good choice for our data. In fact, this behavior can be even more drastically observed when attempting to use the QMC-Voronoi algorithm to train the network. It is noteworthy that all points from the set $\mathcal{X}$ are assigned to only two points (in fact, the first two points) in the set $P$. Consequently, these points are the only points for which a corresponding response exists and, therefore, the only points to train the neural network with. Given that two points are insufficient for training a neural network, this method could not be included in Table (6).
The reason why $K$-means clustering is not performing well may be attributed to the clustering based on the $\boldsymbol{x}$-space. This causes many clusters to contain points with different corresponding responses. This is not a problem for a cluster like $\{2,2,2,2,2,2,2,4,4\}$, since the rounded cluster mean response is still $2$. Though, clusters such as $\{2,2,2,2,4,4,4,4\}$ correspond to a response of $3$, which is not even a response within the cluster. Therefore incorporating the $y$-space into the clustering method appears to be a logical approach. This could also explain why the robust version performs less effectively in this instance. It incorporates the $\boldsymbol{x}$-space into the loss function instead of just the $y$-space. As a result, for this discrete $y$-space problem, a clustering approach based on the output space, performs the best. One potential solution to this issue is to employ a different rounding method. Instead of taking the average of the responses, we could select the response that appears the most frequently. However, this does not address the fundamental challenge that methods primarily using the $\boldsymbol{x}$-space to cluster can produce clusters with a multitude of corresponding responses.
@@
The supercompress is the only method that predicts every number (with a certain degree of accuracy). The robust version fails to predict zeros. Consequently, it overpredicts other numbers, particularly high numbers such as $7$ and $8$. This is the primary reason for the lower accuracy compared to the normal supercompress. The $K$-means clustering method drastically overestimates $2$ and $3$ to be the correct numbers. This systematic error highlights the ineffectiveness of this method in this context. The most unsatisfactory result is produced by QMC-averaging, which predicts only one number. This is not surprising, given that only two points have notable weights and therefore dominate the loss function. The reason for the absence of certain classes in the prediction can be explained by Table (7), which shows the distribution of the corresponding responses among the original and compressed data sets. It should be noted that due to rounding, the percentages may not sum to $100$. Additionally, there is a greater concentration of mass around high numbers. This is a consequence of the fact that many clusters contain two points. If the average value has a $0.5$ part, the response is rounded up, for example, the response corresponding to the set $\{8,9\}$ is 9.